In a current carrying solenoid, determine the đ‘‘đ”/đ‘‘đ‘„ value at the (i) both ends and (ii) centre of the solenoid.


1
Expert's answer
2022-03-21T12:47:25-0400

We know that

(A)



Both end


Magnetic field End of solenoid

"dB=\frac{\mu_0 NIR^2}{2(R^2+x^2)^\frac{3}{2}}"

"\frac{dB}{dx}=-\frac{\mu_0 ni}{2}\int sin\theta d\theta"

"B=-\frac{\mu_0 ni}{2}\int _{\frac{\pi}{2}}^\pi sin\theta d\theta"

"B=\frac{\mu_0 ni}{2}[cos\theta ]_\frac{\pi}{2}^\pi"

"B=\frac{\mu_0 ni}{2}"

(B)

Magnetic field Centre of solenoid

"B_c=-\frac{\mu_0 ni}{2}\int _{0}^\pi sin\theta d\theta"

"B_c=\frac{\mu_0 ni}{2}[cos\theta ]_0^\pi"

"B_c=\mu_0ni"